OpenStudy (anonymous):

h(t)=cot t, [pi/4,3pi/4] find the avg rate of change of the function over the given intervals HELP!!

OpenStudy (anonymous):

Average rate of change is \[ \frac{h(t_2)-h(t_1)}{t_2-t_1} \]

OpenStudy (anonymous):

In this case \(t_1=\pi /4\) and \(t_2=3\pi/4\)

OpenStudy (anonymous):

I understand that bit I don't get how you can plug the values in and get (-4/pi)

OpenStudy (anonymous):

What is \(t_2-t_1\)?

OpenStudy (anonymous):

pi/2?

OpenStudy (anonymous):

Yep. What is \(\cot(3\pi/4)-\cot(\pi/4)\)

OpenStudy (anonymous):

cot(pi/2)

OpenStudy (anonymous):

Nope! You must evaluate \(\cot\) first!

OpenStudy (anonymous):

What is \(\cot (3\pi/4)\)?

OpenStudy (anonymous):

1.169422 or 1.17

OpenStudy (anonymous):

No, \(\cot(3\pi/4)=-1\)

OpenStudy (anonymous):

And \(\cot(\pi/4)=1\)

OpenStudy (anonymous):

Then simplify\[ \frac{-1-1}{\frac \pi 2} \]

OpenStudy (anonymous):

Okay I see how you got that now, I'm a little concerned though as to why my calculator gives me 1.69 for the \[\cot 3\pi \div4\]

OpenStudy (anonymous):

im in radians

OpenStudy (anonymous):

cot (3 x pi/4)

OpenStudy (anonymous):

Never mind, I figured it out haha. Thanks for your help!!